The hypotenuse of a right triangle is 26 millimeters. One leg of the right triangle is 10 millimeters. What is the length of the other leg?
18 millimeters
20 millimeters
24 millimeters
28 millimeters

Answer by addingup(3677)

Pythagoras says: a^2 + b^2 = c^2, where c^2 is the hypotenuse
Using this formula, your problem says:
a^2 + 10^2 = 26^2
a^2 = 26^2 - 10^2
a^2 = 676 - 100
a^2 = 576
(sqrt a^2) = sqrt 576
a = 24